# MTR++: Multi-Agent Motion Prediction with Symmetric Scene Modeling and Guided Intention Querying

### v1

@@ 1 Introduction
In contrast to the initial MTR framework presented in our preliminary version [46], which focuses on the multimodal motion prediction of a single agent, we introduce an advanced MTR++ framework that extends the capability to predict multimodal motion concurrently for multiple agents (see Fig. 1). Instead of individually encoding the scene context around each agent as in the preliminary version, we propose a novel symmetric scene context modeling strategy. This strategy employs a shared context encoder to symmetrically encode the entire scene for each agent, incorporating a novel query-centric self-attention module to jointly capture the intricate scene context information within their respective local coordinate systems. Furthermore, we introduce mutually-guided intention quering module in the motion decoder network, enabling agents to interact and influence each other’s behavior. This facilitates more precise and scene-compliant joint motion prediction for multiple agents. Through these two enhancements, experimental results demonstrate that compared to the initial MTR framework, the MTR++ framework effectively predicts more accurate multimodal future trajectories for multiple agents simultaneously. Additionally, as shown in Fig. 1, the efficiency advantage of the MTR++ framework becomes more pronounced as the number of agents increases.
@@ 2 Related Work
Simultaneous Motion Prediction of Multiple Agents. In predicting an individual agent’s future trajectories, state-of-the-art works [21, 50], including our preliminary version [46], typically customize the scene context encoding for that agent by normalizing all inputs centered on it. This strategy results in computational inefficiencies when predicting motion for multiple agents. To simultaneously predict future trajectories for multiple agents, SceneTransformer [39] encodes all road graphs and agents into a scene-centric embedding applicable to all agents. However, their feature encoding still relies on a global coordinate system centered on an agent of interest (e.g., the autonomous vehicle), limiting its performance for off-center agents. Recent works [69, 27] explore encoding the agents’ node features in an ego-centric coordinate system, while they generally construct hand-crafted relation graphs and alternate node-edge updating strategy. In contrast, our MTR++ framework introduces symmetric scene context modeling for all agents with innovative query-centric self-attention, operating on a straightforward polyline graph using the native transformer encoder module with relative position encoding, thereby promoting more efficient and concise shared scene context encoding.
@@ 3 MTR for Multimodal Motion Prediction
We propose Motion TRansformer (MTR), which adopts a novel transformer encoder-decoder architecture incorporating iterative motion refinement for predicting multimodal future motion. The overall framework is presented in Figure 2. In Sec. 3.1, we introduce our encoder network for scene context modeling. In Sec. 3.2, we present motion decoder network with a novel concept of intention query for predicting multimodal trajectories. Finally, in Sec. 3.3, we introduce the optimization process of our framework.
@@ 3.2 Motion Decoder with Intention Query
Learnable Intention Query111 To streamline the illustration of the motion decoder, we simplify the two components of the motion query pair in our preliminary version [46] by using the new concept of intention query.. To efficiently and precisely pinpoint an agent’s potential motion intentions, we propose the learnable intention query to diminish the uncertainty of future trajectories by employing different intention queries for different motion modes. Specifically, for each category, we generate $\mathcal{K}$ representative intention points $I^{(\text{s})}\in\mathbb{R}^{\mathcal{K}\times 2}$ (where “s” indicating a single focal agent) by utilizing the k-means clustering algorithm on the endpoints of ground-truth (GT) trajectories in the training dataset (refer to Fig. 3). Each intention point embodies an implicit motion mode, accounting for both motion direction and velocity. Given the intention points of a single focal agent, we model each intention query as the learnable positional embedding of a specific intention point:
@@ 3.3 Multimodal Prediction with Gaussian Mixture Model
Training Loss. Given the predicted Gaussian Mixture Models for a specific future time step, we adopt negative log-likelihood loss to maximum the likelihood of the agent’s ground-truth position $(\hat{Y}_{x},\hat{Y}_{y})$ at this time step, and the detailed loss can be formulated as:
@@ 5.2 Main Results
Performance comparison for joint motion prediction. We also evaluate the proposed MTR frameworks on the joint motion prediction benchmark, merging the marginal predictions of two interacting agents into a joint prediction as explicated in [7, 15, 47]. We select the top 6 joint predictions from 36 potential combinations of these agents, with the confidence of each combination being the product of marginal probabilities. As indicated in Table II, our initial MTR framework already surpasses state-of-the-art approaches [39, 47] by substantial margins on all measures, reducing the miss rate from $49.42\%$ to $44.11\%$ and enhancing the mAP from $12.39\%$ to $20.37\%$. Furthermore, our advanced MTR++ framework, which allows us to concurrently predict future motion for two interactive agents with shared context encoding, amplifies the robust performance of MTR across all metrics, achieving an increase of $+2.89\%$ in terms of mAP and reducing the miss rate by $2.68\%$. The extraordinary performance enhancements of the MTR++ framework emphasize that, through the adoption of symmetric scene context encoding and mutually-guided intention querying, our MTR++ framework can accurately predict future trajectories that exhibit scene consistency among highly interacting agents. Additionally, we also provide some qualitative results in Fig. 9 to show our predictions in complicated interacting scenarios. Notably, as of May 30, 2023, our MTR++ framework holds the $1^{st}$ rank on the joint motion prediction leaderboard of WOMD [56],
@@
Performance comparison on the Argoverse 2 dataset. As shown in Table III, we also provide the performance comparison of our approach on the Argoverse 2 dataset for reference. Since our latest MTR++ framework is designed for simultaneous motion prediction of multiple focal agents, we only evaluate our MTR framework on this dataset, which follows the common marginal motion prediction setting of this dataset. We compare our approach with the top-10 submissions on the leaderboard of Argoverse 2 dataset [1] at the time of our MTR framework submission. These submissions, primarily developed for the Argoverse 2 Motion Forecasting Competition 2022, represent highly competitive approaches. Notably, our MTR framework achieves new state-of-the-art performance with remarkable improvements in miss-rate-related metrics, highlighting the exceptional generalizability and robustness of our approach.
@@
TABLE III: The performance comparison on the test set leaderboard of the Argoverse 2 dataset. $K$ is the number of predicted trajectories for calculating the evaluation metrics.
@@ 5.3 Ablation Study
TABLE IV: Effects of different strategies for generating trajectories from encoded context features in the MTR framework.
Effects of the learnable intention query. We investigate the effectiveness of different strategies for generating future trajectories based on encoded context features. These strategies include the simple MLP head [27, 39], the goal-based head [21], the head with 6 latent anchor embeddings [50], and the head with the learnable intention query. The first four rows of Table IV illustrate the performance comparison of these strategies, where our proposed learnable intention query demonstrates significantly superior results. Specifically, our strategy achieves a much better mAP compared to the previous latent anchor embedding [50] (i.e., +5.53%) and dense-goal-based methods [67, 21] (i.e., +4.67%). This improvement can be attributed to our mode-specific querying strategy, where each intention query is associated with an explicit intention point, enabling more accurate and precise multimodal predictions.
@@
TABLE V: Effects of different strategies for generating intention points.
Effects of the distribution of intention points. As introduced in Sec. 3.2, we utilize the k-means clustering algorithm to generate 64 intention points, which serve as the foundation for our intention queries. In order to compare this approach with the straightforward uniform sampling strategy, we uniformly sample 8×8 = 64 intention points by considering the range of trajectory distribution for each category (see Fig. 3). The results presented in Table V indicate a significant drop in performance when replacing the k-means clustering algorithm with uniform sampling for generating intention points. This comparison highlights the superiority of our k-means clustering algorithm, as it produces a more accurate and comprehensive distribution of intention points. Consequently, it effectively captures the diverse future motion intentions of our interested agent with a small number of intention points.
Effects of the iterative trajectory refinement. In Sec. 3.1, we introduce the utilization of stacked transformer decoder layers for iterative refinement of predicted trajectories by continually aggregating fine-grained features via dynamic map collection. As shown in the last two rows of Table IV, this iterative refinement approach significantly reduces the miss rate metric by 1.48% and improves the performance of mAP by +1.6%. By continually aggregating trajectory-specific features from the context encoder with the proposed intention queries, the refinement process effectively improves the accuracy and quality of the predicted trajectories.
@@
TABLE VI: Effects of local self-attention in the transformer encoder of the MTR framework. “#polyline” is the number of input map polylines used for context encoding. “OOM” indicates running out of memory.
Effects of local attention for context encoding. Table VI demonstrates that the utilization of local self-attention in our context encoders leads to slightly superior performance compared to global attention when using the same number of map polylines as input. This finding confirms the significance of incorporating the input’s local structure for more effective context encoding, and the inclusion of such prior knowledge through local attention positively impacts performance. Moreover, local attention proves to be more memory-efficient, allowing for performance improvements even when increasing the number of map polylines from 256 to 1,024. In contrast, global attention suffers from memory limitations due to its quadratic complexity.
@@
TABLE VII: Effects of the symmetric scene context encoder and mutual-guided intention query in the MTR++ framework. The performance evaluation is conducted using a maximum of 8 given focal agents from the validation set of WOMD. To further assess the efficiency of different models as the number of focal agents increases, we set 8, 16, and 32 fake agents as the focal agents for calculating their inference efficiency and memory efficiency.
Effects of the symmetric scene context modeling module. In Sec. 4.1, we present the symmetric scene context encoding module, which utilizes a shared context encoder for motion prediction of multiple interested agents in the same scene. Table VII demonstrates the effectiveness of incorporating our symmetric context encoder into the MTR framework (referred to as MTR+). With MTR+, we achieve comparable performance to the MTR framework while significantly reducing both inference latency and memory cost. Specifically, when the number of interested agents increases from 8 to 32, MTR requires individual scene context encoding for each agent, causing a substantial increase in inference latency and memory cost. In contrast, MTR+ utilizes the query-centric self-attention module to encode the entire scene with a shared symmetric context encoder, leading to a remarkable reduction in inference latency (from 193ms to 98ms for 32 interested agents) and memory cost (from 15.6GB to 4.7GB for 32 interested agents). Furthermore, we provide a breakdown analysis of inference latency in Figure 7, which demonstrates that as the number of interested agents increases, the latency of MTR’s context encoder significantly rises, while the latency of MTR+’s context encoder remains constant due to the utilization of shared context features, since these shared context features enable the prediction of future trajectories for any number of agents within the scene.
@@
Effects of the mutually-guided intention querying strategy. Building upon our proposed symmetric scene context encoder for joint motion prediction, we introduce the mutually-guided intention querying strategy in Sec.4.2. This strategy enables the interaction of future behaviors among multiple agents through the propagation of information among their intention queries. In TableVII, we observe that the mutually-guided intention querying strategy significantly enhances the performance of MTR+ with a remarkable mAP improvement of $+2.49\%$. This improvement demonstrates the effectiveness of broadcasting the potential future behaviors of each agent to other agents via their intention queries, allowing MTR++ to predict more confident future behaviors by considering the overall development of the scene elements.
Furthermore, as each agent incorporates multiple intention queries (i.e., 64 in MTR frameworks), we investigate the interaction among these intention queries within each agent and across different agents. As presented in Table VIII, removing either type of interaction results in a significant decrease in performance by at least $-2.13\%$ in terms of mAP. Removing both types of interaction leads to a larger performance drop of $-2.70\%$ in terms of mAP. This analysis highlights the importance of the interaction among an agent’s different intention queries, enabling the generation of more accurate multimodal future trajectories. Additionally, the interaction among intention queries across different agents empowers the model to predict informed and scene-compliant future trajectories for multiple agents, thereby yielding additional performance improvement.
@@
TABLE VIII: Effects of the interaction of intention queries in the MTR++ framework. “Within Each Agent” and “Across Different Agents” indicates enabling the information interaction within the intention queries of each agent and across the intention queries of different agents, respectively.
Effects of the query-centric self-attention. We introduce the query-centric self-attention module in Sec. 4.1, which plays a vital role in modeling the relationship between tokens within their respective local coordinate systems. This module enables both symmetric scene context modeling and mutually-guided intention querying. In Table IX, we examine the effects of different positional encoding strategies in query-centric self-attention. The results in the first three rows indicate that query-centric relative positional encoding is crucial for achieving optimal performance. Removing this encoding or replacing it with global positional encoding significantly decreases performance by $-2.41\%$ and $-2.91\%$ in terms of mAP, respectively. This finding demonstrates the importance of modeling the relationship in the local coordinate system of each query token, as it benefits the simultaneous motion prediction for multiple agents by treating all tokens symmetrically. Additionally, comparing the last three rows of Table IX, we observe that adding positional embeddings to both the query/key tokens and value tokens yields the best performance.
@@
TABLE IX: Effects of the query-centric self-attention module in the MTR++ framework. “With Query/Key” and “With Value” indicates adding the position embedding to query/key tokens and value tokens, respectively.
@@
TABLE X: Effects of the dense future prediction module in the MTR++ framework.
@@
TABLE XI: Effects of the number of decoder layers in the MTR++ framework.
Effects of dense future prediction. We investigate the impact of the dense future prediction module in Table X. By removing this module, we observe a significant decrease in the performance of the MTR++ framework, with a -1.48% drop in mAP. We attribute this result to the beneficial effects of the dense future prediction module. It not only provides dense supervision for the context encoder, enabling it to learn more effective features for motion prediction of all agents in the scene, but also enhances the motion decoding process in the decoder network by incorporating agent features with their potential future trajectories, thereby enriching the contextual information for multimodal motion prediction.
Effects of the number of decoder layers. We investigate the number of transformer decoder layers in the MTR++ framework in Table XI. We observe a consistent improvement in performance as we increase the number of decoder layers from 1 to 6. This improvement can be attributed to the stacked transformer decoder layers with the mutually-guided intention querying module, which facilitates the generation of more scene-compliant future trajectories through iterative trajectory refinement based on the predicted behaviors of other agents. However, increasing the number of decoder layers to 9 does not yield further improvement, suggesting a diminishing return. As a result, we adopt 6 decoder layers in our MTR++ framework to strike a balance between performance and efficiency.

### v2

@@ 1 Introduction
In contrast to the initial MTR framework presented in our previous version [46], which focuses on the multimodal motion prediction of a single agent, we introduce an advanced MTR++ framework that extends the capability to predict multimodal motion concurrently for multiple agents (see Fig. 1). Instead of individually encoding the scene context around each agent as in the previous version, we propose a novel symmetric scene context modeling strategy. This strategy employs a shared context encoder to symmetrically encode the entire scene for each agent, incorporating a novel query-centric self-attention module to jointly capture the intricate scene context information within their respective local coordinate systems. Furthermore, we introduce mutually-guided intention quering module in the motion decoder network, enabling agents to interact and influence each other’s behavior. This facilitates more precise and scene-compliant joint motion prediction for multiple agents. Through these two enhancements, experimental results demonstrate that compared to the initial MTR framework, the MTR++ framework effectively predicts more accurate multimodal future trajectories for multiple agents simultaneously. Additionally, as shown in Fig. 1, the efficiency advantage of the MTR++ framework becomes more pronounced as the number of agents increases.
@@ 2 Related Work
Furthermore, we employ Gaussian Mixture Models in tandem with intention queries to model the continuous distribution of an agent’s multimodal future behavior at each time step. This approach yields a parametric multimodal future distribution as the output, capable of generating occurrence probabilities for any specified future trajectory. In contrast, utilizing a set of sparse trajectories fails to provide a continuous and compact future distribution. Similarly, predicting dense future heatmaps incurs a substantial computational cost, necessitating a compromise between resolution and computational efficiency.
Simultaneous Motion Prediction of Multiple Agents. In predicting an individual agent’s future trajectories, state-of-the-art works [21, 50], including our previous version [46], typically customize the scene context encoding for that agent by normalizing all inputs centered on it. This strategy results in computational inefficiencies when predicting motion for multiple agents. To simultaneously predict future trajectories for multiple agents, SceneTransformer [39] encodes all road graphs and agents into a scene-centric embedding applicable to all agents. However, their feature encoding still relies on a global coordinate system centered on an agent of interest (e.g., the autonomous vehicle), limiting its performance for off-center agents. Recent works [69, 27] explore encoding the agents’ node features in an ego-centric coordinate system, while they generally construct hand-crafted relation graphs and alternate node-edge updating strategy. In contrast, our MTR++ framework introduces symmetric scene context modeling for all agents with innovative query-centric self-attention, operating on a straightforward polyline graph using the native transformer encoder module with relative position encoding, thereby promoting more efficient and concise shared scene context encoding.
@@ 3 MTR for Multimodal Motion Prediction
We propose Motion TRansformer (MTR), which adopts a novel transformer encoder-decoder architecture incorporating iterative motion refinement for predicting multimodal future motion. The overall framework is presented in Fig. 2. In Sec. 3.1, we introduce our encoder network for scene context modeling. In Sec. 3.2, we present motion decoder network with a novel concept of intention query for predicting multimodal trajectories. Finally, in Sec. 3.3, we introduce the optimization process of our framework.
@@ 3.2 Motion Decoder with Intention Query
Learnable Intention Query111 To streamline the illustration of the motion decoder, we simplify the two components of the motion query pair in our previous version [46] by using the new concept of intention query.. To efficiently and precisely pinpoint an agent’s potential motion intentions, we propose the learnable intention query to diminish the uncertainty of future trajectories by employing different intention queries for different motion modes. Specifically, for each category, we generate $\mathcal{K}$ representative intention points $I^{(\text{s})}\in\mathbb{R}^{\mathcal{K}\times 2}$ (where “s” indicating a single focal agent) by utilizing the k-means clustering algorithm on the endpoints of ground-truth (GT) trajectories in the training dataset (refer to Fig. 3). Each intention point embodies an implicit motion mode, accounting for both motion direction and velocity. Given the intention points of a single focal agent, we model each intention query as the learnable positional embedding of a specific intention point:
@@ 3.3 Multimodal Prediction with Gaussian Mixture Model
Training Loss. Given the predicted Gaussian Mixture Models for a specific future time step, we adopt negative log-likelihood loss to maximize the likelihood of the agent’s ground-truth position $(\hat{Y}_{x},\hat{Y}_{y})$ at this time step, and the detailed loss can be formulated as:
@@ 5.2 Main Results
Method
Number of
Parameters
Inference
Latency
Miss Rate $\downarrow$ mAP normal-$\uparrow$
${}^{\dagger}$ceneTransformer [39] 15.3M 52ms (V100) 0.1564 0.2788
DenseTNT [21] 1.1M 540ms 0.1573 0.3281
HDGT [27] 12.1M 1320ms 0.1511 0.2854
MTR++ (light) 11.7M 67ms 0.1430 0.3896
MTR 65.8M 193ms 0.1351 0.4129
MTR++ 86.6M 118ms 0.1298 0.4329
TABLE III: Performance is compared in terms of the number of parameters and inference latency for multi-agent prediction, utilizing the Waymo Open Motion Dataset. Latency is estimated by mandating each model to predict the motion of 32 agents per scenario, leveraging an NVIDIA Quadro RTX 8000 GPU. $\dagger$: the number of parameters and latency are reported in the respective source paper, using NVIDIA V100 GPU.
@@
Performance comparison for joint motion prediction. We also evaluate the proposed MTR frameworks on the joint motion prediction benchmark, merging the marginal predictions of two interacting agents into a joint prediction as explained in [7, 15, 47]. We select the top 6 joint predictions from 36 potential combinations of these agents, with the confidence of each combination being the product of marginal probabilities. As indicated in Table II, our initial MTR framework already surpasses state-of-the-art approaches [39, 47] by substantial margins on all measures, reducing the miss rate from $49.42\%$ to $44.11\%$ and enhancing the mAP from $12.39\%$ to $20.37\%$. Furthermore, our advanced MTR++ framework, which allows us to concurrently predict future motion for two interactive agents with shared context encoding, amplifies the robust performance of MTR across all metrics, achieving an increase of $+2.89\%$ in terms of mAP and reducing the miss rate by $2.68\%$. The extraordinary performance enhancements of the MTR++ framework emphasize that, through the adoption of symmetric scene context encoding and mutually-guided intention querying, our MTR++ framework can accurately predict future trajectories that exhibit scene consistency among highly interacting agents. Additionally, we also provide some qualitative results in Fig. 7 to show our predictions in complicated interacting scenarios. Notably, as of May 30, 2023, our MTR++ framework holds the $1^{st}$ rank on the joint motion prediction leaderboard of WOMD [56].
Moreover, we provide the performance comparison in terms of the number of parameters and inference latency in Table III. To enable a comparison of performance and latency with a similar parameter count, we have provided a streamlined version of MTR++ (notated as “MTR++ (light)") by diminishing the number of encoder layers to 2 and decoder layers to 3. As indicated in Table III, our lightweight MTR++ surpasses existing approaches markedly, enhancing the mAP from $32.81\%$ to $38.96\%$. When compared with SceneTransformer [39], our MTR++ (light) not only employs fewer parameters ($15.3$M down to $11.7$M) but also attains superior performance ($27.88\%$ boosting to $38.96\%$). When compared with DenseTNT [21], albeit our MTR++ (light) utilizes more parameters, it operates considerably faster (67ms versus 540ms). This enhanced speed can be attributed to our method sidestepping the densely sampled goal points present in DenseTNT. When compared with HDGT [27], our MTR++ (light) runs significantly faster (67ms as opposed to 1320ms). This efficiency arises from our usage of the well-optimized native Transformer module, whereas HDGT deploys a custom graph neural network. Furthermore, our standard MTR and MTR++ elevate performance through the use of additional parameters, while simultaneously maintaining inference latencies significantly lower than the existing open-source approaches [21, 27].
@@
Performance comparison on the Argoverse 2 dataset. As shown in Table IV, we also provide the performance comparison of our approach on the Argoverse 2 dataset for reference. We compare our approach with the top-10 submissions on the leaderboard of Argoverse 2 dataset [1] at the time of our MTR framework submission. These submissions, primarily developed for the Argoverse 2 Motion Forecasting Competition 2022, represent highly competitive approaches. Notably, our MTR framework achieves new state-of-the-art performance with remarkable improvements in miss-rate-related metrics. Moreover, our MTR++ further surpasses the performance of both MTR and other existing approaches across all three metrics, thereby highlighting the exceptional generalizability and robustness of our approach.
@@
TABLE IV: The performance comparison on the test set leaderboard of the Argoverse 2 dataset. $K$ is the number of predicted trajectories for calculating the evaluation metrics.
@@ 5.3 Ablation Study
TABLE V: Effects of different strategies for generating trajectories from encoded context features in the MTR framework.
Effects of the learnable intention query. We investigate the effectiveness of different strategies for generating future trajectories based on encoded context features. These strategies include the simple MLP head [27, 39], the goal-based head [21], the head with 6 latent anchor embeddings [50], and the head with the learnable intention query. The first four rows of Table V illustrate the performance comparison of these strategies, where our proposed learnable intention query demonstrates significantly superior results. Specifically, our strategy achieves a much better mAP compared to the previous latent anchor embedding [50] (i.e., +5.53%) and dense-goal-based methods [67, 21] (i.e., +4.67%). This improvement can be attributed to our mode-specific querying strategy, where each intention query is associated with an explicit intention point, enabling more accurate and precise multimodal predictions.
@@
TABLE VI: Effects of different strategies for generating intention points.
@@
TABLE VII: Effects of local self-attention in the transformer encoder of the MTR framework. “#polyline” is the number of input map polylines used for context encoding. “OOM” indicates running out of memory.
Effects of the distribution of intention points. As introduced in Sec. 3.2, we utilize the k-means clustering algorithm to generate 64 intention points, which serve as the foundation for our intention queries. In order to compare this approach with the straightforward uniform sampling strategy, we uniformly sample 8×8 = 64 intention points by considering the range of trajectory distribution for each category (see Fig. 3). The results presented in Table VI indicate a significant drop in performance when replacing the k-means clustering algorithm with uniform sampling for generating intention points. This comparison highlights the superiority of our k-means clustering algorithm, as it produces a more accurate and comprehensive distribution of intention points. Consequently, it effectively captures the diverse future motion intentions of our interested agent with a small number of intention points.
Effects of the iterative trajectory refinement. In Sec. 3.1, we introduce the utilization of stacked transformer decoder layers for iterative refinement of predicted trajectories by continually aggregating fine-grained features via dynamic map collection. As shown in the last two rows of Table V, this iterative refinement approach significantly reduces the miss rate metric by 1.48% and improves the performance of mAP by +1.6%. By continually aggregating trajectory-specific features from the context encoder with the proposed intention queries, the refinement process effectively improves the accuracy and quality of the predicted trajectories.
@@
TABLE VIII: Effects of the symmetric scene context encoder and mutual-guided intention query in the MTR++ framework. The performance evaluation is conducted using a maximum of 8 given focal agents from the validation set of WOMD. To further assess the efficiency of different models as the number of focal agents increases, we set 8, 16, and 32 fake agents as the focal agents for calculating their inference efficiency and memory efficiency.
Effects of local attention for context encoding. Table VII demonstrates that the utilization of local self-attention in our context encoders leads to slightly superior performance compared to global attention when using the same number of map polylines as input. This finding confirms the significance of incorporating the input’s local structure for more effective context encoding, and the inclusion of such prior knowledge through local attention positively impacts performance. Moreover, local attention proves to be more memory-efficient, allowing for performance improvements even when increasing the number of map polylines from 256 to 1,024. In contrast, global attention suffers from memory limitations due to its quadratic complexity.
Effects of the symmetric scene context modeling module. In Sec. 4.1, we present the symmetric scene context encoding module, which utilizes a shared context encoder for motion prediction of multiple interested agents in the same scene. Table VIII demonstrates the effectiveness of incorporating our symmetric context encoder into the MTR framework (referred to as MTR+). With MTR+, we achieve comparable performance to the MTR framework while significantly reducing both inference latency and memory cost. Specifically, when the number of interested agents increases from 8 to 32, MTR requires individual scene context encoding for each agent, causing a substantial increase in inference latency and memory cost. In contrast, MTR+ utilizes the query-centric self-attention module to encode the entire scene with a shared symmetric context encoder, leading to a remarkable reduction in inference latency (from 193ms to 98ms for 32 interested agents) and memory cost (from 15.6GB to 4.7GB for 32 interested agents). Furthermore, we provide a breakdown analysis of inference latency in Fig. 8, which demonstrates that as the number of interested agents increases, the latency of MTR’s context encoder significantly rises, while the latency of MTR+’s context encoder remains constant due to the utilization of shared context features, since these shared context features enable the prediction of future trajectories for any number of agents within the scene.
@@
Effects of the mutually-guided intention querying strategy. Building upon our proposed symmetric scene context encoder for joint motion prediction, we introduce the mutually-guided intention querying strategy in Sec.4.2. This strategy enables the interaction of future behaviors among multiple agents through the propagation of information among their intention queries. In TableVIII, we observe that the mutually-guided intention querying strategy significantly enhances the performance of MTR+ with a remarkable mAP improvement of $+2.49\%$. This improvement demonstrates the effectiveness of broadcasting the potential future behaviors of each agent to other agents via their intention queries, allowing MTR++ to predict more confident future behaviors by considering the overall development of the scene elements.
Furthermore, as each agent incorporates multiple intention queries (i.e., 64 in MTR frameworks), we investigate the interaction among these intention queries within each agent and across different agents. As presented in Table IX, removing either type of interaction results in a significant decrease in performance by at least $-2.13\%$ in terms of mAP. Removing both types of interaction leads to a larger performance drop of $-2.70\%$ in terms of mAP. This analysis highlights the importance of the interaction among an agent’s different intention queries, enabling the generation of more accurate multimodal future trajectories. Additionally, the interaction among intention queries across different agents empowers the model to predict informed and scene-compliant future trajectories for multiple agents, thereby yielding additional performance improvement.
@@
TABLE IX: Effects of the interaction of intention queries in the MTR++ framework. “Within Each Agent” and “Across Different Agents” indicates enabling the information interaction within the intention queries of each agent and across the intention queries of different agents, respectively.
Effects of the query-centric self-attention. We introduce the query-centric self-attention module in Sec. 4.1, which plays a vital role in modeling the relationship between tokens within their respective local coordinate systems. This module enables both symmetric scene context modeling and mutually-guided intention querying. In Table X, we examine the effects of different positional encoding strategies in query-centric self-attention. The results in the first three rows indicate that query-centric relative positional encoding is crucial for achieving optimal performance. Removing this encoding or replacing it with global positional encoding significantly decreases performance by $-2.41\%$ and $-2.91\%$ in terms of mAP, respectively. This finding demonstrates the importance of modeling the relationship in the local coordinate system of each query token, as it benefits the simultaneous motion prediction for multiple agents by treating all tokens symmetrically. Additionally, comparing the last three rows of Table X, we observe that adding positional embeddings to both the query/key tokens and value tokens yields the best performance.
@@
TABLE X: Effects of the query-centric self-attention module in the MTR++ framework. “With Query/Key” and “With Value” indicates adding the position embedding to query/key tokens and value tokens, respectively.
@@
TABLE XI: Effects of the dense future prediction module in the MTR++ framework.
@@
TABLE XII: Effects of the number of decoder layers in the MTR++ framework.
Figure 10: The rank distribution of the best mode’s score of our MTR++ framework on the Waymo Open Motion Dataset.
Figure 11: Qualitative results of the failure cases of MTR frameworks on WOMD. The green curves indicate the ground-truth trajectories of our interested agents (green rectangle). We use gradient color to visualize the predicted trajectory waypoints at different future time steps, and trajectory confidence is visualized by setting different transparent. Abbreviation: Vehicle (V).
Effects of dense future prediction. We investigate the impact of the dense future prediction module in Table XI. By removing this module, we observe a significant decrease in the performance of the MTR++ framework, with a -1.48% drop in mAP. We attribute this result to the beneficial effects of the dense future prediction module. It not only provides dense supervision for the context encoder, enabling it to learn more effective features for motion prediction of all agents in the scene, but also enhances the motion decoding process in the decoder network by incorporating agent features with their potential future trajectories, thereby enriching the contextual information for multimodal motion prediction.
Effects of the number of decoder layers. We investigate the number of transformer decoder layers in the MTR++ framework in Table XII. We observe a consistent improvement in performance as we increase the number of decoder layers from 1 to 6. This improvement can be attributed to the stacked transformer decoder layers with the mutually-guided intention querying module, which facilitates the generation of more scene-compliant future trajectories through iterative trajectory refinement based on the predicted behaviors of other agents. However, increasing the number of decoder layers to 9 does not yield further improvement, suggesting a diminishing return. As a result, we adopt 6 decoder layers in our MTR++ framework to strike a balance between performance and efficiency.
5.4 Discussion of Failure Cases and Future Challenges
While MTR/MTR++ frameworks have demonstrated remarkable performance and achieved a state-of-the-art position, they continue to confront two primary challenges as below.
Quality score estimation: As evident in Table I, our model achieves a relatively lower average precision (i.e., below 0.5), while maintaining a favorable miss rate (approximately 0.13). This discrepancy arises due to the inadequacy of current quality score estimations in accurately reflecting the quality of each predicted trajectory, resulting in significant penalties imposed by the average precision metric. A compelling illustration of this issue is provided in Fig. 10, where our analysis reveals that only 45.5% of agents adhere to the predicted trajectory with the highest quality score. This underscores the limitations of existing quality score predictions, ultimately contributing to suboptimal average precision. Consequently, one important direction for future research involves investigating methods to refine the alignment between predicted scores and the quality of the corresponding trajectory. This would yield a more precise distribution of an agent’s future behaviors, enhancing the predictive capabilities of our models.
Diverse and accurate multimodal behaviors in rare scenarios: Another significant challenge centers on generating diverse and accurate multimodal behaviors that faithfully capture the real intentions of the agent, particularly in rare scenarios. As illustrated in Fig. 11, our model, in certain scenarios, may generate homogenized future trajectories that fail to encompass the genuine intentions of the agent, resulting in imprecise multimodal predictions. Thus, a pivotal future research endeavor involves the development of methods that yield comprehensive multimodal behaviors, ensuring that multiple future trajectories not only exist but also represent distinct, meaningful intentions of the agent rather than merely variations along the same direction.
These challenges and future research directions are instrumental in advancing the capabilities of motion prediction models, and addressing them will contribute to more accurate and comprehensive predictions, ultimately benefitting a wide range of applications.